5. 4.56 QUESTION 24 In a one-way ANOVA there are three independent samples, with n 8 n2 10 and n 7; the calculated F-statistic is F 3.95; at the 0.05 level of significance, what conclusion would be reached? 1. Fail to reject Ho. 2. Reject Ho 3. The population means are the same. 4. Need data set. 5. None of the above 25
Answer #1

for above degree of freedom of numerator =k-1=3-1=2

degree of freedom for denominator =n1+n2+n3-k=8+10+7-3=22

for 0.05 level and (2,22) degree of freedom ; critical value =3.44

as test statistic 3.95 is higher then critical value ; we reject null hypothesis

option 2 is correct , Reject Ho